Indicate which of the following nuclear equations are correct: 3.9 Pu – 350 + He - 339 cºc-75 Ru + 4n
Answer #1

11 99 Tc Ru tie 43 3:44. left hand side (2015) Atomic number : 43 Mass number = qq Right hand side (R.A.s) Atomic number = 44257 (iv) 249 cc 12 + 7 ... Rut to 4 98 n 104 LHS Atomic number = 98 +6 = 104 Mass number = 249 +12=261 RiMOS Atomic number